# RadGPT: Constructing 3D Image-Text Tumor Datasets

### v1

@@ Abstract
With over 85 million CT scans performed annually in the United States, creating tumor-related reports is a challenging and time-consuming task for radiologists. To address this need, we present RadGPT, an Anatomy-Aware Vision-Language AI Agent for generating detailed reports from CT scans. RadGPT first segments tumors, including benign cysts and malignant tumors, and their surrounding anatomical structures, then transforms this information into both structured reports and narrative reports. These reports provide tumor size, shape, location, attenuation, volume, and interactions with surrounding blood vessels and organs. Extensive evaluation on unseen hospitals shows that RadGPT can produce accurate reports, with high sensitivity/specificity for small tumor ($<$2 cm) detection: 80/73% for liver tumors, 92/78% for kidney tumors, and 77/77% for pancreatic tumors. For large tumors, sensitivity ranges from 89% to 97%. The results significantly surpass the state-of-the-art in abdominal CT report generation.
RadGPT generated reports for 17 public datasets. Through radiologist review and refinement, we have ensured the reports’ accuracy, and created the first publicly available image-text 3D medical dataset, comprising over 1.8 million text tokens and 2.7 million images from 9,262 CT scans, including 2,947 tumor scans/reports of 8,562 tumor instances. Our reports can: (1) localize tumors in eight liver sub-segments and three pancreatic sub-segments annotated per-voxel; (2) determine pancreatic tumor stage (T1-T4) in 260 reports; and (3) present individual analyses of multiple tumors—rare in human-made reports. Importantly, 948 of the reports are for early-stage tumors.
Figure 1: AbdomenAtlas 3.0 is a large-scale, image-text tumor dataset of 9,262 3D CT scans. Each CT scan has per-voxel tumor annotations and reports, including 3,036 liver tumors, 354 pancreatic tumors and 4,239 kidney tumor, 5,262 of which are small tumor ($\leq$2cm). In addition, AbdomenAtlas 3.0 provides detailed annotations for pancreatic cancer staging (T1–T4), as well as per-voxel segmentations of liver sub-segments (1–8) and pancreatic sub-segments (head, body, and tail).
@@ 1 Introduction
Each year, over 85 million CT scans are performed in the United States [62, 55], growing 6% per year, and significantly outpacing the 0.7% annual growth rate of the medical imaging workforce [16]. This disparity puts radiologists under significant time pressure, making it challenging to generate detailed, accurate radiology reports. These reports involve identifying tumors, measuring them, determining their location, classifying types [^1] We use the term ‘tumor’ to refer to an abnormal mass of tissue, which may be malignant (cancer) or benign (e.g., cysts)., and assessing surrounding tissue interactions—a process that is labor-intensive, time-consuming, and prone to human error. Recent studies showed that radiologists have a sensitivity rate of about 65–85% for detecting small liver tumors [52] and as low as 33–44% for small pancreatic tumors [24], leading to frequent oversight. These missed detections can delay timely treatment and diminish the quality of AI training, as many small tumors go unrecorded in radiology reports. Automated report generation is needed to improve the accuracy and efficiency of tumor detection beyond human-made reports.
Automated report generation from abdominal CT scans is particularly important as many high-mortality cancer like liver, pancreatic, and colorectal cancer originate in the abdomen and require early detection and precise staging [19, 70, 13, 40, 41]. However, as of this writing, there are no publicly available datasets or models for automated report generation that detail tumor-related findings using abdominal CT scans [^2] Two concurrent studies are underway, but the datasets and models from Merlin [9] are not publicly available, and M3D [4] provides CT scans in a 2D PNG/JPEG format (0–255), unsuitable for clinical use (§2). Pre-existing report generation approaches, such as those based on Vision-Language Models (VLMs), are mainly designed for 2D images [68, 35, 12, 54, 59, 67] or Chest CT scans [65, 27, 66, 5], using different imaging protocols from abdominal scans and relying on weak annotations from CLIP-like [50] image-text pairs, which limits their ability to provide precise location, measurement, and tracking for individual tumors (evidenced in Table 2).
To address this gap, we present Radiology Generative Pre-trained Transformers, RadGPT, an Anatomy-Aware Vision-Language AI Agent for generating detailed reports from abdominal CT scans. This approach seeks to improve the accuracy of tumor quantification and bring consistency to report generation. RadGPT can convert per-voxel annotation of tumors and organs into two types of reports: (1) structured reports and (2) narrative reports. As shown in Figure 1, the structured reports follow a radiologist-provided template, and RadGPT can adapt this reports into narrative reports that match the style (word choice and text organization) of a target institution (hospital or clinic). Both formats have been validated by radiologists and offer comprehensive and detailed information, such as tumor size, location, attenuation, volume, descriptions of adjacent organs (e.g., fatty liver, enlarged spleen), patient demographics, contrast enhancement protocols, and pancreatic tumor T stage, which consider tumor size and interaction with blood vessels, key considerations for surgical removal. In addition to generating accurate reports, the proposed RadGPT offers two key functions:
  1. 1.
    Enhancing human-made reports: RadGPT can leverage per-voxel tumor segmentation and human-made reports or clinical notes (i.e., informal and short reports) to generate fusion reports3.3). They inherit the precision and detail of segmentation (e.g., multiple precise tumor size and volume measurements), and the diagnostic breadth of human-made reports/notes, encompassing multiple conditions beyond tumors (§4.5). This synergy enhances clarity and enriches human-made reports/notes with quantitative insights.
  2. 2.
    Diagnostic evaluation of AI-made reports: RadGPT introduces a quantitative diagnostic assessment for report generation—a feature missing in current approaches for abdominal CT report generation (§2). First, a LLM extracts numerical and categorical labels (e.g., tumor presence/absence) from AI- and human-made reports, with 96% accuracy in zero-shot (Figure 3). By comparing the labels from AI- and human-made reports, we can evaluate AI’s diagnostic accuracy using metrics such as sensitivity and specificity (§3.4).
Finally, we used RadGPT to generate reports for 17 public datasets, and through annotation, review, and refinement by 12 radiologists, we created AbdomenAtlas 3.0 (Figure 1), the largest public image-text abdominal CT dataset. It contains 1,843,262 text tokens from reports and 2,789,975 images from 9,262 CT scans. Each scan features structured and narrative reports detailing tumors’ sizes, shapes, locations and appearances. The dataset includes a total of 2,947 reports describing 8,562 tumor instances across the liver (929 reports, 3,036 instances), pancreas (344 reports, 354 instances), and kidneys (1,674 reports, 4,239 instances), all per-voxel annotated. Additionally, 6,061 reports document no tumors in these organs, serving as control. Notably, a large number of reports documenting tumors document small tumors ($\leq$2 cm)—347 (37.4%) for liver, 83 (24.1%) for pancreas, and 466 (27.8%) for kidney—critical for early cancer detection but rare and challenging to collect. Tumors are localized using per-voxel annotated liver sub-segments (1–8) and pancreatic sub-segments (head, body, tail). AbdomenAtlas 3.0 includes 260 reports with pancreatic tumor staging (T1–T4), enabled by per-voxel annotations of key vessels (e.g., SMA). AbdomenAtlas 3.0 is the first public dataset with pancreatic cancer staging and pancreas sub-segments. It also provides 240 human-AI fusion reports, combining our structured/narrative reports with clinical notes (see §3.3). Metadata like contrast enhancement, slice thickness, and in-plane spacing further enhance AbdomenAtlas 3.0.
@@
FLARE’23 [2022] [link] 4,100 35 1 0 $\rightarrow$ 376 0 $\rightarrow$ 34 0 $\rightarrow$ 763
@@
CT-ORG [2020] [link] 140 8 6 0 $\rightarrow$ 41 0 0 $\rightarrow$ 17
Trauma Det. [2023] [link] 4,714 23 13 0 $\rightarrow$ 46 0 $\rightarrow$ 15 0 $\rightarrow$ 38
@@
AbdomenCT-1K [2021] [link] 1,050 12 7 0 $\rightarrow$ 82 0 $\rightarrow$ 100 0 $\rightarrow$ 113
MSD CT Tasks (6) [2021] [link] 945 1 1 251 191 0 $\rightarrow$ 224
WORD [2021] [link] 120 1 1 0 $\rightarrow$ 28 0 0 $\rightarrow$ 26
AMOS [2022] [link] 200 2 1 0 $\rightarrow$ 54 0 $\rightarrow$ 3 0 $\rightarrow$ 41
AbdomenAtlas 3.0 (ours) 9,262 88 19 301 $\rightarrow$ 929 191 $\rightarrow$ 344 452 $\rightarrow$ 1,674
@@
Table 1: Besides being the only public abdominal CT dataset with paired radiology reports, AbdomenAtlas 3.0 offers 3.12$\times$ more annotated tumors then the combined total of its constituent datasets. The table highlights how AbdomenAtlas 3.0 enhances public datasets with reports and tumor annotations. It includes 936 CT scans with liver tumors, 342 with pancreatic tumors, and 1,416 with kidney tumors, most newly annotated with radiologist support. Each sample includes per-voxel annotations and reports. AbdomenAtlas 3.0 is also the first dataset to provide per-voxel segmentations of pancreas sub-segments and peripancreatic blood vessels.
@@
Figure 2: RadGPT adopts 3-stage pipeline for report generation. Blue arrows denote quality assurance steps used for creating AbdomenAtlas 3.0 but omitted in fully-automated generation. Stage I. Segmentation. We utilize DiffTumor [14] and a proprietary-trained nnU-Net [29] to segment 26 anatomical structures important for cancer detection and staging3. For AbdomenAtlas 3.0, we asked radiologists to revise segmentations, or substituted them by the ground-truth from public datasets when available. Stage II. Structured Report Generation. Deterministic algorithms (§3.1.1-3.1.3) extract radiologist-selected attributes (important for cancer detection, staging and treatment planning) from CT scans and per-voxel segmentations to fill a template created by radiologists, generating structured reports for liver, kidney, and pancreatic cancer assessment. The rule-based deterministic algorithms ensure the reports are fully coherent with the segmentations and explainable. Stage III. Style Adaptation. A LLM adapts structured reports to a target hospital’s narrative style using in-context learning with example reports from the hospital. An new LLM-based example selection prioritizes reports with similar diagnoses (§3.2). The LLM is instructed to maintain medical information accuracy and perform double checks for consistency. For AbdomenAtlas 3.0, radiologists conducted final report revisions. Additionally, we use an LLM to fuse structured reports with human-made reports, creating comprehensive fusion reports that combine segmentation-based precision with the generality of human-made reports (Section 3.3).
@@ 2 Related Work
Per-voxel tumor annotations are scarce. Public abdominal CT datasets typically focus on one tumor type—such as liver [8], pancreas [3], and kidney [25] tumors—and present only a few hundreds per-voxel tumor annotations (Table 1). This limited quantity undermines robust AI training and evaluation. Thus, supported by radiologists, we tripled the number of per-voxel tumor annotations in the 17 public datasets inside AbdomenAtlas 3.0 (Table 1).
Even scarcer than per-voxel tumor annotations are reports. No public abdominal CT dataset has real-world radiology reports (i.e., from a medical institution). Actually, only one public abdominal CT dataset, M3D-Cap [4], has some text annotations. They are captions, sourced from the Radiopedia website [22], along with CT scans. However, these scans do not meet clinical standards. They are 2D JPG/PNG image series instead of standard 3D NIfTI or DICOM volumes. Thus, they lack information that is essential for precise tumor measurement by AI, like inter-image spacing and Hounsfield units (HU) [70]. Conversely, the CT scans in AbdomenAtlas 3.0 have this information, being sourced from 88 medical institutions in standard format.
Due to the scarcity of abdominal CT reports in public datasets, only two models for abdominal CT report generation exist: M3D [4] (public) and Merlin [9] (unreleased). However, both were solely evaluated with text similarity metrics, like BLEU and ROUGE [39]. These standard LLM evaluation metrics can be highly influenced by changes in the report style, even when underlying diagnoses do not change (§4.4). In contrast, RadGPT enables evaluating AI-made reports using diagnostic sensitivity and specificity, which are clinically relevant metrics [64, 10].
@@
Table 1 shows advantages of AbdomenAtlas 3.0 over the 17 public datasets. It is the only dataset with reports, and has three report types: structured reports (§3.1), narrative reports (§3.2), and fusion reports (§3.3). All reports were generated by RadGPT, an anatomy-aware vision-language AI agent illustrated in Figure 2. First, RadGPT transformed the ground-truth per-voxel annotations of tumors and organs into structured reports, using deterministic algorithms. These algorithms ensure the diagnoses in the report align with the ground-truth per-voxel annotations. Then, RadGPT used LLMs to convert the structured reports into narrative reports, and to create fusion reports, which fuse structured reports with clinical notes.
@@ 3.1 Creating Structured Reports
Structured reports use a radiologist-designed template, enhancing clarity and aiding medical decisions [1] (Figure 1). To fill out this template, RadGPT performs key tasks: (1) Sub-segments organs to localize tumors (§3.1.1). (2) Measures tumor size, volume, and attenuation from segmentations (§3.1.2). (3) Uses tumor and blood vessel segmentations for staging (§3.1.3).
3.1.1 Sub-segment Organs to Localize Tumors
@@
Human-made reports use organ sub-segments to locate tumors. Location is key for prognosis, tracking tumor progression, and treatment planning. E.g., the possibility of tumor removal depends on its location [57]. To locate liver and pancreas tumors in structured reports, RadGPT sub-segments the organs and checks which sub-segments intersect with the tumor. Besides releasing the first public dataset with pancreas sub-segments, we release the first public AI models that sub-segment pancreas and liver.
@@
For liver sub-segmentation, we propose using ground-truth per-voxel annotations of the entire liver to help the AI find liver sub-segments. First, we offset the liver intensity (by 200 HU), following its ground-truth per-voxel annotation. Second, using these CT scans with offsets as input, we trained an nnU-Net [29] for liver sub-segmentation. The sub-segments follow the Couinaud standard [18], which divides the liver into eight sub-segments that can be independently removed in surgeries. Couinaud sub-segment annotations are publicly available for 131 LiTS CT scans [69, 8], which we used for training. Given the small size of this dataset, we fine-tuned an nnU-Net pre-trained on 9,262 CT scans in AbdomenAtlas 1.1 [36]. After training, we inferenced the nnU-Net on AbdomenAtlas 3.0. The HU value offsetting ensured the precise alignment between the generated sub-segments and the existing liver ground-truth per-voxel annotations. AbdomenAtlas 3.0 is the second [69] but the largest public dataset with liver sub-segments.
@@
For pancreas sub-segmentation, there is no public dataset with per-voxel annotations of pancreas head, body, and tail. Thus, to sub-segment the pancreas, we used a radiology landmark: the superior mesenteric artery (SMA). We trained an nnU-Net to segment the SMA (using private data), and developed a deterministic algorithm that uses the segmented SMA to sub-segment the pancreas (algorithm 1). The algorithm follows standard radiology guidelines. First, it uses the SMA to find the pancreatic neck, since it curves around the SMA. The neck locates the head-body boundary. Then, the body-tail boundary is set at the midpoint along their length [58]. AbdomenAtlas 3.0 is the first public dataset with pancreas sub-segment annotations.
3.1.2 Segment Tumors and Measure Like Radiologists

### v2

@@ Abstract
Cancers identified in CT scans are usually accompanied by detailed radiology reports, but publicly available CT datasets often lack these essential reports. This absence limits their usefulness for developing accurate report generation AI. To address this gap, we present AbdomenAtlas 3.0, the first public, high-quality abdominal CT dataset with detailed, expert-reviewed radiology reports. All reports are paired with per-voxel masks and they describe liver, kidney and pancreatic tumors. AbdomenAtlas 3.0 has 9,262 triplets of CT, mask and report—3,955 with tumors. These CT scans come from 17 public datasets. Besides creating the reports for these datasets, we expanded their number of tumor masks by 4.2×, identifying 3,011 new tumor cases. Notably, the reports in AbdomenAtlas 3.0 are more standardized, and generated faster than traditional human-made reports. They provide details like tumor size, location, attenuation and surgical resectability. These reports were created by 12 board-certified radiologists using our proposed RadGPT, a novel framework that converted radiologist-revised tumor segmentation masks into structured and narrative reports. Besides being a dataset creation tool, RadGPT can also become a fully-automatic, segmentation-assisted report generation method. We benchmarked this method and 5 state-of-the-art report generation vision-language models. Our results show that segmentation strongly improves tumor detection in AI-made reports.
Figure 1: AbdomenAtlas 3.0 is a large-scale, image-text tumor dataset of 9,262 3D CT scans. Each CT scan has per-voxel tumor annotations and reports, including 5,582 liver tumors, 368 pancreatic tumors and 4,424 kidney tumor, 7,003 of which are small tumors ($\leq$2cm). In addition, AbdomenAtlas 3.0 provides detailed annotations for pancreatic cancer staging (T1–T4), as well as per-voxel segmentation of liver sub-segments (1–8) and pancreatic sub-segments (head, body, and tail). Structured, narrative, and enhanced reports were created by a team of 12 board-certified radiologists assisted by our proposed Radiology Generative Pretrained Transformer (RadGPT).
@@ 1 Introduction
Each year, over 85 million CT scans are performed in the United States [52, 44], growing 6% per year, and significantly outpacing the 0.7% annual growth rate of the medical imaging workforce [13]. This disparity puts radiologists under significant time pressure, making it challenging to generate detailed, accurate radiology reports. AI may support report generation, but it requires data. To address this gap, we present AbdomenAtlas 3.0 (summarized in Figure 1 and Table 1), the first high-quality abdominal CT dataset with reports. It has 9,262 3D CTs in NIfTI format (2,789,975 CT slices) sourced from 17 public datasets (Table 1), which originally had no radiology report. 12 board-certified radiologists, assisted by RadGPT (introduced below), generated reports for all CTs—totaling 1,843,262 tokens. For each CT, we document tumor size, location, attenuation (HU), and volume for each identified tumor. Reports also include T-stage for pancreatic cancer (PDAC), derived from tumor size and vessel involvement, critical for surgery. Each CT has both structured (template-based) and narrative (free-text) reports, and precise voxel-level annotations. Reports cover tumors in the liver, pancreas, and kidneys, including 3,011 tumors newly identified by the radiologists. Our reports also describe organ abnormalities (e.g., fatty liver, enlarged spleen), patient demographics, and contrast phase. They locate tumors in liver segments (1–8) and pancreas segments (head, body, tail)—all annotated per-voxel. This is the largest liver sub-segment dataset, and the first public pancreas sub-segment dataset. Also, we enhanced 240 existing human-made reports, covering 66 distinct diagnoses, with more detailed tumor analyses.
@@
FLARE’23 [2022] [link] 4,100 35 1 0 $\rightarrow$ 564 0 $\rightarrow$ 38 0 $\rightarrow$ 941
@@
CT-ORG [2020] [link] 140 8 6 0 $\rightarrow$ 44 0 0 $\rightarrow$ 21
Trauma Det. [2023] [link] 4,714 23 13 0 $\rightarrow$ 113 0 $\rightarrow$ 32 0 $\rightarrow$ 38
@@
AbdomenCT-1K [2021] [link] 1,050 12 7 0 $\rightarrow$ 117 0 $\rightarrow$ 94 0 $\rightarrow$ 181
MSD CT Tasks (6) [2021] [link] 945 1 1 251 $\rightarrow$ 462 191 0 $\rightarrow$ 388
WORD [2021] [link] 120 1 1 0 $\rightarrow$ 47 0 $\rightarrow$ 1 0 $\rightarrow$ 45
AMOS [2022] [link] 200 2 1 0 $\rightarrow$ 74 0 $\rightarrow$ 4 0 $\rightarrow$ 56
AbdomenAtlas 3.0 (ours) 9,262 138 19 301 $\rightarrow$ 1,472 191 $\rightarrow$ 361 452 $\rightarrow$ 2,122
@@
[…]
To create AbdomenAtlas 3.0, we developed Radiology Generative Pre-trained Transformers (RadGPT), an anatomy‑aware vision‑language AI agent that assists radiologists in creating CT-report datasets. We started with our previous AbdomenAtlas 1.1 [31], composed of 17 public datasets and their organ segmentation masks, but no tumor segmentation nor report. First, RadGPT segments liver, kidney, and pancreas tumors, along with liver/pancreas sub-segments, surrounding organs, and blood vessels1. Then, radiologists revise the segmented tumors, annotating missed ones and removing false positives. We call the dataset with CT scans and tumor segmentation masks AbdomenAtlas 2.0, and we also release it here. From the revised segmentations, RadGPT extracts attributes (e.g., tumor size, volume, attenuation, stage) via deterministic, rule-based algorithms. These attributes are used to fill a radiologist-designed template, producing structured reports. RadGPT’s deterministic algorithms ensure that the structured reports are fully explainable and fully coherent with the radiologist-revised segmentations. Next, RadGPT converts the structured reports into free-text narrative reports, using large language models (LLMs) that emulate the style (word choice and organization) of radiologists at a major US hospital—through in-context learning with special example selection (§3.2). Last, RadGPT fuses per-voxel segmentations with human-made reports/clinical notes to produce enhanced human reports3.3), combining precise and detailed tumor analysis from segmentation with broader diagnostic range (66 diagnoses) from human-made reports. Reports were verified by radiologists (Appendix C). We call the final triplet dataset—CT scans, tumor masks, reports—AbdomenAtlas 3.0.
We evaluated six CT report generation models on AbdomenAtlas 3.0 (internal validation) and a private dataset (external validation): CT2Rep [21], M3D [4], CT-CHAT [20], Merlin [8], RadFM [54] and RadGPT. Besides a dataset creation tool, RadGPT can also become a fully-automatic, segmentation-assisted report generation model, by converting the outputs of a segmentation model into reports, without radiologist revision. We expect AbdomenAtlas 3.0 to foster segmentation-assisted report generation, as the dataset has CTs, per-voxel annotations and reports. We evaluated all report generation models with a new diagnostic metric (§B.3). It first uses an LLM to extract labels (tumor presence) from AI- and human-made reports. Then, it compares the labels from AI- and human-made reports to evaluate AI’s diagnostic sensitivity and specificity (§3.4). To validate this new metric, radiologists manually evaluated LLM labeling—it achieved 96% zero-shot accuracy (Figure 4). Our contributions are:
  1. 1.
    AbdomenAtlas 3.0 is the first public dataset with high-quality abdominal CT scans (9,262), radiology reports (structured, narrative, and enhanced), and tumor masks.
  2. 2.
    With 12 radiologists, we annotated 3,011 new tumors in the 17 public datasets inside AbdomenAtlas 3.0—expanding their number of tumor masks by 4.2×.
  3. 3.
    Our reports locate liver and pancreas tumors within sub-segments of the organs. They also measure contact between tumors and blood vessels for pancreatic tumor staging. Staging and sub-segments are key for surgery.
  4. 4.
    We developed Rad-GPT to assist dataset creation: unlike current VLMs, it uses deterministic algorithms to convert radiologist-revised tumor masks into reports, improving reports’ trustworthiness and interpretability. Also, RadGPT can generate fully-automated reports.
  5. 5.
    We benchmarked 5 SOTA VLMs for report generation and showed segmentation improves report generation.
@@
Figure 2: The RadGPT 3-stage pipeline for report generation. Blue arrows denote human revision used to create AbdomenAtlas 3.0. By skipping revision, RadGPT can also become a fully-automatic, segmentation-assisted report generation AI. Stage I. Segmentation. DiffTumor [11] and nnU-Net [25] segment 26 anatomical structures important for cancer detection and staging1. Radiologists corrected wrong tumor segmentations in AbdomenAtlas 3.0, and ground-truths from public datasets were used when available. Stage II. Structured Report Generation. Deterministic algorithms (§3.1.1-3.1.3) extract radiologist-selected attributes—important for cancer detection, staging and treatment—from CTs and segmentations. Attributes fill a radiologist-designed template, generating structured reports detailing liver, kidney, and pancreatic tumors. The rule-based deterministic algorithms ensure the reports are fully coherent with segmentations and explainable. Stage III. Style Adaptation. LLM adapts structured reports into a target hospital’s narrative style, leveraging example reports from the hospital—in-context learning prioritizing examples of similar diagnoses (§3.2). LLM is asked to preserve medical information and double checks for consistency. Radiologists revised reports in AbdomenAtlas 3.0. Also, LLM can fuse structured and human-made reports, creating enhanced human reports combining segmentation-based precision with humans’ broad diagnostic range (§3.3).
@@ 2 Related Work
Per-voxel tumor annotations are scarce. Most public abdominal CT datasets concentrate on a single tumor type (e.g., liver [7], pancreas [3], or kidney [22]) and contain only a few hundred tumor annotations (Table 1). This small volume of annotations hinders effective AI training and evaluation. To address this, our radiologists have quadrupled the number of per-voxel tumor annotations in the 17 public datasets included in AbdomenAtlas 3.0 (Table 1).
Real-world radiology reports are even rarer than per-voxel tumor annotations. At the time of writing, no publicly available abdominal CT dataset contains authentic clinical reports. Only one dataset, M3D-Cap [4], provides textual captions (sourced from Radiopaedia [19]), but its scans are 2D JPG/PNG image series rather than standard 3D NIfTI or DICOM volumes. Consequently, crucial information such as inter-slice spacing and Hounsfield units (HU) is missing [59]. In contrast, CT scans in AbdomenAtlas 3.0 were collected in standard formats from 138 medical institutions, retaining clinically important metadata. As another unique quality, our reports are paired with tumor and organ masks—fostering the development of segmentation-assisted report generation AI.
Due to the scarcity of reports in public datasets, only two models specifically target abdominal CT report generation: M3D [4] (publicly released) and Merlin [8] (partially released). Text-similarity metrics were used to evaluate both models (e.g., BLEU and ROUGE [35]; Merlin was also evaluated with RadGraph-F1), but these metrics can be skewed by style variations even when the underlying diagnoses remain unchangedB.3). In contrast, we propose the evaluation of AI-generated reports using diagnostic sensitivity and specificity (Table 2)—clinically meaningful and acceptable metrics [55, 9]. Lastly, although many report-generation models exist for 2D X-ray [51, 12, 33, 34, 57, 45], adapting them to 3D CT may require profound re-design, which may unfairly represent the originals. Why? First, tumors in CTs can occupy $\leq$0.0001% of the full volume, vs.  5–10% in X-rays. Second, many X-ray models rely on 2D pre-trained models, but CT data is 3D. Processing CT slices individually is computationally prohibitive, and it is difficult to align slices with findings in reports. Thus, all models we evaluated in AbdomenAtlas 3.0 [21, 4, 20, 8, 54] are designed for CT.
@@
[…]
@@ 3.1 Creating Structured Reports
Structured reports use a radiologist-designed template, enhancing clarity and aiding medical decisions [1] (Figure 1). To fill the template, RadGPT uses segmentation and deterministic algorithms to: (1) sub-segment organs to locate tumors (§3.1.1); (2) measure tumor size, volume, and attenuation (§3.1.2); (3) perform cancer staging from tumor and blood vessel segmentations (§3.1.3).
3.1.1 Sub-segment Organs to Locate Tumors
@@
Human-made reports use organ sub-segments to locate tumors. Location is key for prognosis, tracking tumor progression, and treatment planning. E.g., the possibility of tumor surgical removal depends on its location [47]. To locate liver and pancreas tumors in structured reports, RadGPT sub-segments the organs and checks which sub-segments intersect with the tumor. RadGPT segments tumors with DiffTumor [11], a public segmentation model, and radiologists revise the segmentations (Appendix C).
@@
For liver sub-segmentation, we leverage whole-liver ground-truth per-voxel annotations to help the AI find liver sub-segments. First, we offset the liver intensity (by 200 HU), following its ground-truth per-voxel annotation. Second, using these CT scans with offsets as input, we trained an nnU-Net [25] for liver sub-segmentation. The sub-segments follow the Couinaud standard [15], which divides the liver into eight sub-segments that can be independently removed in surgeries. Couinaud sub-segment annotations are publicly available for 131 LiTS CT scans [58, 7], which we used for training. Given the small size of this dataset, we fine-tuned an nnU-Net pre-trained on 9,262 CT scans in AbdomenAtlas 1.1 [31]. After fine-tuning, we inferenced the nnU-Net on AbdomenAtlas 3.0. The HU value offsetting ensured the precise alignment between the generated sub-segments and the existing liver ground-truth per-voxel annotations. AbdomenAtlas 3.0 is the second [58] but largest public dataset with liver sub-segments.
@@
For pancreas sub-segmentation, there is no public dataset with per-voxel annotations of pancreas head, body, and tail—ours is the first. Thus, to subsegment the pancreas, we used the superior mesenteric artery (SMA) as a landmark. We trained an nnU-Net to segment the SMA (using private data) and developed a deterministic algorithm that uses the SMA segmentation to sub-segment the pancreas (Sup. Alg. 1). First, it uses the SMA to find the pancreatic neck, since it curves around the SMA. The neck locates the head-body boundary. Then, the body-tail boundary is set at the midpoint along their length. Our landmark-based deterministic algorithm closely mimics how radiologists use mesenteric vessels to subsegment the pancreas [48]. AbdomenAtlas 3.0 is the first public dataset with pancreas sub-segments.
3.1.2 Measure Tumors Like Radiologists